Determine the mass of air in a vacant room whose dimensions are 4m × 5m × 6m at 100 kPa and 25°C.


1
Expert's answer
2021-07-07T08:39:35-0400

"pV=\frac mMRT,"

"m=\frac{pabcM}{RT}=\frac{10^5\cdot 4\cdot 5\cdot 6\cdot 29\cdot 10^{-3}}{8.31\cdot 298}=140.5~kg."